find two positive numbers whose product is 64 and whose sum is a minimum


1
Expert's answer
2021-05-13T11:55:50-0400

Let the numbers be x and y


According to question-


"xy=64\Rightarrow y =\dfrac{64}{x}~~~~~~~-(1)"


Also Let S be their sum-


"S=x+y

\\[9pt]

S=x+\dfrac{64}{x}"


Differentiate w.r.t. x-


"\dfrac{dS}{dx}=1-\dfrac{64}{x^2}"


Putting "\dfrac{dS}{dx}=0\Rightarrow 1-\dfrac{64}{x^2}=0\Rightarrow x^2=64\Rightarrow x=\pm 8"


Again differentiating "\dfrac{dS}{dx}" w.r.t. "x-"


"\dfrac{d^2S}{dx^2}=\dfrac{128}{x^3}"


"\dfrac{d^2S}{dx^2}_{\text{ at } x=8}=\dfrac{128}{8^3}=2"


So "\dfrac{d^2S}{dx^2}>0"


SO, S is miniumum at x=8


Then Other number "y=\dfrac{64}{8}=8"


Hence The two numbers are 8 and 8.